Find density of an object Measure all the required dimensions of the cylinder using Vernier Caliper and calculate the density of the cylinder Cylinder Number:2_ Mass of Cylinder: 45.2 Measurements: ドイ Value Quantity Height H Diameter l4.3 cm cm Inner Diameter 2.3 cm Inner Height h. cm Relevant Formulas: Volume of a cylinder (radius r, height h) r h Volume of a cylinder (radius r, height h)-πrzh please fine
a) volume of cavity
_________cm^3

b) volume of entire object calculation
_________cm^3

c) density of cylinder calculation
  _________kg/m^3
Answer #1

here,

a)

volume of cavity , Vc = pi * (d/2)^2 * h

Vc = pi * ( 0.023/2)^2 * ( 0.004)

Vc = 1.66 * 10^-6 m^3

Vc = 1.66 cm^3

b)

volume of entire object , V = Volume of cyclinder - volume of cavity

V = pi * (D/2)^2 * H - 1.66 * 10^-6 m^3

V = pi * ( 0.043/2)^2 * 0.014 - 1.66 * 10^-6

V = 1.87 * 10^-5 m^3

V = 18.7 cm^3

c)

mass of cyclinder , m = 45.2 g = 0.0452 kg

density , d = mass /volume

d = 0.0452 /( 1.87 * 10^-5)

d = 2.42 * 10^3 kg/m^3